OpenStudy (anonymous):

integrate (1-x^2)/[x(1-2x)]dx

OpenStudy (anonymous):

\[\int\frac{1-x^2}{x(1-2x)}dx\] Have you tried partial fraction decomposition?

hartnn (hartnn):

partial fraction cannot work, for that degree of numerator must be strictly less than degree of denom.

OpenStudy (anonymous):

So long division first, then PFD.

hartnn (hartnn):

yes.

OpenStudy (anonymous):

I'd suggest rewriting as \[\int\frac{x^2-1}{2x^2-x}dx\] That should help with the long division

OpenStudy (anonymous):

k

hartnn (hartnn):

x^2-1 = 1/2 (2x^2-2) = (1/2) (2x^2-x +x-2)

hartnn (hartnn):

if you wantt o avoid long division....

hartnn (hartnn):

got that ? ^ :P

OpenStudy (anonymous):

hmmm..i have solveed it acc to 2 u @hartnn and i got 1/2[1+(1/2x-1)-(2/2x^2-x)]

hartnn (hartnn):

you got that as final answer ? and where is +c ?

OpenStudy (anonymous):

nahi yaar till now i have only put value

OpenStudy (anonymous):

yup i got the answer and have checked it from wolfram Thank You so much @hartnn

hartnn (hartnn):

ohh...i was verifying manually :P welcome ^_^